Nov 26th to Dec 5th 2007: First scientific run

Daniel Bonneau, Jean-Michel Clausse, Denis Mourard

December 2nd to 5th

  • During these nights, we have been able to record two sets of data mainly for qualification purposes . The seeing conditions were really not good with only two sequences of a few hours with a seeing below 1.5''.

Friday November 30th and Saturday december 1st

  • The bad weather conditions did not permit observations during these nights, except for a short control of the new alignment.
  • We have installed in the lab a light detector for the ALGOLs security. After many attempts it seems to work correctly now and the HV are switched off if the light is turn on in the lab.
  • The observing schedule has been prepared carefully and all the templates are now ready thanks to the VEGA_PLAN software.
  • Many improvement in the control software as well as a lot of time spent on the bias detector problem. The shielding does not improve much the situation.
  • The VPN box between MtWilson and Grasse Observatory begins to work but the configuration has been difficult, thanks to the firewall protections.
  • We are now installed (royalement!) in the Kaypten cottage. Thanks Hal!

Thursday November 29th

  • New alignment of VEGA/CHARA: the M2&M6 are restored to their initial positions. The sky pupil is reinstalled on its nominal position. RM is aligned again on the sky pupil. Then a lateral shift appears on the four beams on the close targets (the same as yesterday night in fact). This shift is first roughly compensated by moving all the M6 mirrors. Then the classical procedure is operated with the close and far targets. After that, the beam samplers are set in position and the alignement of the VEGA beams is made on the CHARA tool on the OPLE rails. Small adjustement of the M6 are necessary. Then the corner cubes near the CHARA-BRT are installed and we can follow the return beam at the entrance of the tip/tilt sensor and at the entrance slit of VEGA. On the slit viewer the return beams appear slightly on the edge of the slit.
  • In depth analysis of the communication of the control software to lighten the internet transfer.
  • The polarisation measurements have been completed by installing the white light source just in front of one BCP and directed towards the spectrograph. Analysis in progress.
  • We also prepare carefully the strategies of the observations, combining the use of W1,W2, E1 and E2 but by pairs only.
  • We have been able to check our new alignment on the sky during the holes in the clouds but seeing was horrible (r0 about 2cm) and no fringes were possible. Forecast for the coming nights are really not good with rain and snow!

Wednesday November 28th

  • First step of the measurements for testing the internal polarisation of VEGA and CHARA. We use the white light source installed before the beam splitters and directly on the CAU. No major differences are detected but new measures are necessary to avoid the reflexion on the CSM beam splitter in the CAU configuration.
  • Continuation of the work on the detectors. The variance of the dark on ALGOLR is 5 ADU for a mean value of 70 ADU. It should be noted also that a slope of about 6 ADU occurs in the vertical direction of the dark (integration of the rows) whereas an important modulation (almost 10%) from column to column appears in the horizontal direction (integration of the columns). These characteristics of the detectors are then found again when working in photon counting regime and it explains why we have this strong modulation on the spectrum.
  • Continuation of the work on the control software.
  • Observations in parallel to MIRC on HR1903 (eps Ori). We have a problem in the alignment: our four beams (E1,S1,W1,W2) are shifted horizontally at the entrance of the spectrograph by about 1 to 2mm. The problems comes from a bad procedure in the alignment made yesterday (probleme with the pupil mask). For continuing the nights we decided to center the beams and the pupils by moving the M2 and M6 mirrors.
  • It should be noted that the four pupils look pretty better than in spetember.
  • Fringes were tracked by MIRC and as soon as we opened the detectors, fringes were present on VEGA with a small offset of 750µm. Observations were made on baseline W1-W2 in HRS mode of VEGA. During this night, three combiners have worked in parallel (CLASSIC, MIRC and VEGA): a great milestone for CHARA!

 Tuesday November 27th

  • The VEGA computers are now correctly installed in the control room.
  • The new version of the detector’s software has been tested. Darks with or without the high voltage have been compared and no differences are detectable. Thus darks will be recorded with the high voltages OFF. Raw images have been recorded for the accurate determination of the low and high threshold. For both detectors, values have been found to be 6 (low) and 180 (high).
  • The alignment of the Image and Pupil Sensor has been corrected.
  • The fine rotation of the detectors has been also adjusted.
  • The alignment of VEGA to CHARA has been made completely, starting from the pupil mask, the M6 and then the BCPs. The beams from the BCP have been carefully aligned with the M1-M2 to the CHARA targets but due to the residual defaults of the BCP some minor motions on the M6 are necessary. After the final alignement, references have been taken at the output of the BCP. These positions as well as the position of the beams at the output of the M6 have to be checked carefully before moving again the M1-M2. Keeping these positions is necessary to avoid changes in the internal optical path between M1-M2.
  • Bad weather for the night…




November 2007: Preparation of the first science run of VEGA on CHARA

The team currently prepares the coming run. Some technical works are under progress, such as:

  • a passive security system to detect light in the optical lab for the security of Vega intensified detectors,
  • a new chiller for Vega detectors that should arrive mid November,
  • the connection of Vega computers and electronics on the new UPS installed by Nils,
  • and the correct installation of VEGA computers in the control room.

Concerning the observations, VEGA is planned on 4 nights: Dec. 2nd to 5th.
>From Nov, 27th to Dec, 1st, we plan to verify optical alignment, install new softwares, continue tests and adjustments on the detectors. Parallel observations with MIRC should also allow us to find our OPD offsets and adjust them.

During a short stay on the mountain (nov. 3th to 5th), intensive new software tests have been made as well as technical works on the detectors to solve the current issues on bad photon detection. See report here (in french).

Stay tuned to follow the mission diary         



See September  2007 integration mission news on this page

See May  2007 news on this page

See April  2007 news on this page

See February & March 2007 news on this page

See January 2007 news on this page

See December 2006 news on this page

See July to November 2006 news on this page

July 2006 moving spectrograph from Calern/GI2T to Grasse/Optical lab

Vega in SPIE 2006 see publications page

Chara agreement April 2006